# arXiv:2603.01198v1[eess.SY] 01 Mar 2026

### v1

@@ 4.5 Limitations and future work
Several limitations should be noted. First, the cooling tower fan model employs a simplified power-approach relationship calibrated to DOE performance curves, rather than a manufacturer-specific model for the Frontier installation. The CT fan savings should therefore be interpreted as directionally correct rather than exact. Second, the optimization assumes that the supply temperature setpoint can be independently adjusted from the cooling tower operation, which may require control system modifications at the facility level. Third, the current study uses a static (point-by-point) optimization with retrospective ramp-rate enforcement; a full receding-horizon MPC formulation with multi-step look-ahead capability could further improve performance during rapid load transients. Finally, the digital twin validation (ASHRAE Guideline 14 compliant, as reported in Section 3.1) was performed using a steady-state energy balance with an adopted $c_{p}$ that differs by 6% from the dataset’s internal conversion constant (Section 2.3.2); transient validation during load ramps and independent calorimetric measurement of $c_{p}$ would strengthen confidence in the absolute predictions.
@@ 5 Conclusion
This paper developed a digital twin for the Frontier exascale supercomputer’s liquid cooling system and established three progressively complex optimization strategies through one full year of operational data. The study produced four principal findings.
First, the dominant source of cooling energy consumption is not where conventional intuition suggests. Cooling tower fan energy accounts for 73% of the total cooling energy budget at Frontier, while pump energy accounts for only 27%. This asymmetry means that optimizing pump flow in isolation, despite the attractive cubic relationship between flow and pump power, captures only a fraction of the available savings. Strategy A (flow-only optimization) achieves 20.4% total savings, but Strategy B (co-optimization of flow and supply temperature) reaches 30.1% by reducing the cooling tower workload through a modest mean supply temperature increase of 2.7C. The co-optimized strategy uses more pump energy than the flow-only strategy, yet achieves greater total savings. This counterintuitive result underscores the importance of system-level thinking over component-level optimization, a principle with broad applicability across coupled thermal systems.
Second, the gap between theoretical optimality and practical implementability is surprisingly small for this system. Strategy C, which enforces ramp-rate constraints of $\pm$50 kg/s per step and $\pm$1C per step on the co-optimized solution, recovers 92.4% of the unconstrained savings (27.8% vs. 30.1%). The implementability cost of 2.3% is modest in absolute terms and does not vary substantially across seasons (1.5–2.6%). This finding suggests that for large-scale liquid cooling systems with smooth thermal dynamics, practitioners can pursue near-optimal energy performance without sacrificing the actuator smoothness and operational predictability that real-world deployment demands.
@@
Third, the proposed methodology is robust to parametric uncertainty. Across the full range of plausible pump power exponents ($n=2$ to $n=3$), Strategy C achieves 21–28% total cooling energy savings. Even under the most conservative assumption ($P\propto\dot{m}^{2}$), the savings exceed 21%, providing a reliable lower bound for facility managers who may lack detailed pump curve data for installed equipment.
@@
Fourth, the digital twin validation demonstrates that a physics-based model with a single calibration parameter (glycol mass fraction) can predict per-subloop return temperatures to within 1.96–2.67% CV-RMSE against ASHRAE Guideline 14 criteria, using the Modelica Buildings Library as the modeling foundation. This confirms the viability of equation-based digital twins as optimization testbeds for HPC cooling infrastructure, where the physical transparency of the model is essential for operator trust and regulatory acceptance.
From a methodological standpoint, this work introduces two concepts that extend beyond the specific application. The implementability gap metric provides a formal, reproducible measure of the energy penalty imposed by practical constraints on theoretically optimal control. The layered optimization framework, which progresses from single-variable to multi-variable to constrained formulations, offers a systematic protocol for decomposing the sources of energy savings and evaluating the cost of operational realism. These concepts are applicable to any coupled thermal system where the gap between simulation-derived optima and deployable control strategies is a concern.
The computational efficiency of the SLSQP-based optimizer (under 10 ms per timestep) makes online supervisory control feasible within a 10-minute update cycle, and the Modelica framework enables straightforward adaptation to alternative cooling architectures through component substitution. As HPC facilities continue to scale toward multi-exascale power envelopes, the gap between theoretical and implementable efficiency will only widen unless it is systematically measured. The tools introduced here provide a foundation for that measurement.

### v2

@@ 4.4 Limitations and future work
Several limitations should be noted. The cooling tower fan model employs a simplified power-approach relationship rather than a manufacturer-specific model, so CT fan savings should be interpreted as directionally correct rather than exact. The optimization assumes that the supply temperature setpoint can be independently adjusted from cooling tower operation, which may require control system modifications. The current study uses static point-by-point optimization with retrospective ramp-rate enforcement; a full receding-horizon MPC formulation with multi-step look-ahead could further improve performance during rapid load transients. The digital twin validation was performed using a steady-state energy balance with an adopted $c_{p}$ that differs by 6% from the dataset’s internal conversion constant; transient validation during load ramps and independent calorimetric measurement of $c_{p}$ would strengthen confidence in the absolute predictions.
@@ 5 Conclusion
This paper developed a Modelica-based digital twin for the Frontier exascale supercomputer’s liquid cooling system and evaluated three progressively constrained optimization strategies across one full year of operational data. The study produced four principal findings.
First, cooling tower fan energy accounts for 73% of the total cooling energy budget, making supply temperature co-optimization essential. Flow-only optimization (Strategy A) achieves 20.4% total savings, but co-optimization of flow and supply temperature (Strategy B) reaches 30.1% through a mean supply temperature increase of 2.7C. The co-optimized strategy uses more pump energy yet achieves greater total savings, underscoring the importance of system-level optimization over component-level approaches.
Second, the implementability gap is surprisingly small. Strategy C, enforcing ramp-rate constraints of $\pm 50$ kg/s per step and $\pm 1$ C per step, recovers 92.4% of the unconstrained savings (27.8% vs. 30.1%). This demonstrates that near-optimal energy performance is achievable without sacrificing actuator smoothness.
@@
Third, the methodology is robust to parametric uncertainty: across pump power exponents $n=2$ to $n=3$, Strategy C achieves 21–28% total savings, providing a reliable bound for facilities lacking detailed pump curve data.
@@
Fourth, a physics-based digital twin with a single calibration parameter (glycol mass fraction) predicts per-subloop return temperatures to within 1.96–2.67% CV-RMSE, confirming the viability of equation-based digital twins as optimization testbeds for HPC cooling infrastructure.
The implementability gap metric and layered optimization framework introduced here are applicable to any coupled thermal system where the cost of practical constraints on theoretically optimal control must be quantified. As HPC facilities scale toward multi-exascale power envelopes, systematic measurement of this gap will become increasingly important.
@@ 7 Data and code availability
Appendix A Supplementary Material
Appendix A
A.1 Data center cooling system optimization: approaches and limitations
Table A1: Comparison of representative cooling system optimization studies. Abbreviations: DT = digital twin; HPC = high-performance computing; DC = data center.
Study Method Application DT Constraints Gap metric
Evans and Gao (2016) Neural net Google DC No No No
Lazic et al. (2018) MPC Google DC No Partial No
Li et al. (2020) Deep RL Enterprise DC No No No
Sarkar et al. (2024) Multi-agent RL Enterprise DC Yes No No
Zhan et al. (2025) Offline RL Production DC No Partial No
Zhu and Lin (2024) Real-time opt. Enterprise DC Yes Partial No
Jadhav and Liu (2026) ML surrogate Frontier (HPC) No No No
Brewer et al. (2024) — (V&V only) Frontier (HPC) Modelica No
Fu et al. (2019a) — (Sim. only) Air-cooled DC Modelica No
This study SLSQP (layered) Frontier (HPC) Modelica Ramp-rate Yes
A.2 Monthly savings heatmap by strategy
Figure A1 provides a compact month-by-month view of cooling-energy savings across all optimization strategies. The heatmap makes the seasonal structure more explicit than the monthly bar chart in the main text. Strategy A contributes relatively modest savings during winter (approximately 6–10%), when the baseline operating point is already close to thermally efficient conditions and the available margin for flow reduction is limited. Its contribution rises substantially during summer, reaching about 25–37%, when baseline flow is most inflated relative to thermal need.
Strategies B and C consistently outperform Strategy A across the year because co-optimization exploits both flow reduction and supply-temperature reset. Their largest absolute advantage over Strategy A appears during the shoulder seasons (March–April and October–November), where setpoint adjustment remains effective even when incremental pump savings are smaller than in summer. The difference between Strategies B and C is smallest during summer, typically around 1–3%, indicating that the ramp-rate constraints impose the least penalty precisely when the achievable savings are greatest.
Overall, the appendix heatmap complements Figure 6 by showing that the annual savings reported in the main text are not driven by a few isolated months, but instead reflect a consistent seasonal pattern in which co-optimization provides benefits throughout the year while remaining most effective during high-load warm-weather operation.
Figure A1: Monthly total cooling energy savings (%) by control policy. Color intensity indicates magnitude. Co-optimization strategies (B and C) outperform flow-only optimization (A) across all months, with the advantage most pronounced during shoulder and winter seasons.
A.3 Operating envelope shift
Figure A2 presents the joint distribution of flow rate and supply temperature for Baseline, Strategy B, and Strategy C, visualized through kernel density contours overlaid on the operating points. Under baseline operation, the facility occupies a broader envelope centered at 257 kg/s mean flow and 20.1C mean supply temperature, with a bimodal flow structure consistent with the seasonal step change discussed in the main text.
Strategy B shifts the operating envelope toward lower flow rates and higher supply temperatures, centering at 187 kg/s and 22.8C. Strategy C exhibits a nearly identical centroid at 189 kg/s and 22.6C, but with a modestly broader spread, reflecting the smoothing effect of ramp-rate constraints. The close similarity between Strategies B and C confirms that ramp constraints primarily affect transition dynamics rather than the steady-state operating target.
Figure A2: Joint distribution of flow rate and supply temperature for (a) Baseline, (b) Strategy B, and (c) Strategy C. Contour lines show kernel density estimates; dotted lines indicate mean values. Co-optimization shifts the operating envelope toward lower flow and higher supply temperature, while Strategy C closely tracks the Strategy B centroid with modestly broader spread due to ramp-rate constraints.
A.4 Ramp-rate compliance
Figure A3 evaluates whether the optimization commands remain compatible with conservative actuator and process limits. The unconstrained Strategy B produces large step-to-step flow and supply-temperature changes that frequently exceed the adopted limits of $\pm 50$ kg/s and $\pm 1$ C per 10-minute interval. By contrast, Strategy C satisfies both limits by construction, indicating that the ramp-constrained policy remains practically implementable while retaining most of the achievable energy benefit.
Figure A3: Step-to-step ramp-rate comparison between unconstrained Strategy B and ramp-constrained Strategy C during a summer week: (a) flow rate changes with $\pm$50 kg/s limit, and (b) $T_{\mathrm{supply}}$ setpoint changes with $\pm$1C limit. Strategy C respects both operational constraints while Strategy B frequently violates them.
A.5 Pump power duration curves
Figure A4 shows the annual pump power duration curves for the baseline and all optimized strategies. All optimization strategies shift the curve downward relative to the baseline, confirming substantial reduction in pump power across operating hours. The largest reduction is achieved by Strategy A, whereas Strategies B and C retain slightly higher pump power because they trade part of the pump savings for lower cooling-tower fan energy.
Figure A4: Annual pump power duration curves showing the distribution of pump power across all operating hours. All optimization strategies substantially reduce pump power relative to the baseline, with the greatest reductions during high-load summer operation.
A.6 Sensitivity to pump curve exponent
Figure A5 shows the sensitivity of total cooling-energy savings to the assumed pump curve exponent for all three optimization strategies. Savings increase monotonically with exponent across all strategies, consistent with the nonlinear dependence of pump power on flow rate. Co-optimization strategies (B and C) consistently outperform flow-only optimization (A) across the full range of plausible exponents, with the relative advantage most pronounced at lower exponents where cooling-tower fan savings account for a larger share of the total benefit.
Figure A5: Sensitivity of total cooling energy savings to pump curve exponent for all three optimization strategies. Co-optimization (Strategies B and C) consistently outperforms flow-only optimization across all exponents.